• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 11:00:23 on March 6,2009

    Entry number 264871

    This entry is a followup to: 264870

    keyword=Compton orbit lock was not on

    MCC says the Compton orbit lock was not on. It is now. Based on the set value they had (-0.97) it looks like we've drifted quite a bit to get to -1.34.

    I'll be keeping an eye on the rates.

    I spoke to Sirish and a drift on the order of .1 (that is, 100 microns) is not expected to dramatically affect the rates, which is why we didn't see a degradation in photon rate over the time scale we were looking at.